A car initially traveled 35 km due south and then traveled 65 km to the west. What is the car’s resultant displacement?

Scale: 1 cm = 10 km



1
Expert's answer
2021-08-31T17:33:15-0400
"s=\sqrt{35^2+65^2}=73.8\ km\\\theta=\arctan{\frac{35}{65}}=28.3\degree\text{ S of W}"